Explain the THREE (3) models of consumption theory and what distinguishes them in terms of marginal...

  1. Explain the THREE (3) models of consumption theory and what distinguishes them in terms of marginal propensity to consume (MPC).
Answer #1

Theory of consumption was initially propounded by J.M keynes,who stated that consumption pattern of individuals depend upon their income.Thereafter further developments took place to amend the keynesian absolute theory of consumption .These are ---

1)Relative income theory of consumption----:This theory is propounded by Dusenberry. According to this theory ,the consumption expenditure of a person doesn't depends upon his income in relation to others' income. This theory suggests that individuals or households try to imitate the consumption pattern of other families or their neighbours.

As per this theory, there is a non proportional relationship between income and consumption. In the short run MPC < APC. In the long run MPC>APC.

2) Lifecycle theory of consumption -------: This theory propounded by Modigliani, who states that a person plans his consumption expenditure during his lifetime and this consumption expenditure depends upon his current income as well as his future expected income. So, a consumer's lifetime income consists of initial wealth and lifetime earnings. The consumption behaviour can be expressed as C= aplha×W + beta×Y,

where C stands for consumption behaviour,

alpha×W stands for MPC out of Wealth,

Beta×Y stands for MPC out of Income.

The MPC out of lifetime income changes with age.

3) Permanent income theory of consumption ----------: This theory is propounded by Friedman. It is somewhat similar to life cycle hypothesis. The economist here states that consumption expenditure of a person depends upon his long term expected income (permanent income), not upon current level of income.

This theory states that an individual's income has 2 components--->

1) permanent income

2) transitory income

As per this theory, a person's permanent income consists of his long term earnings from employment, retirement pensions and income derived from possessions of capital assets.

Transitory income is a short term temporary income earned through windfall gains, say, winning from lottery.

According to Friedman, consumption is proportional to permanent income as people tend to save their transitory income.

C^p =KY^p

Where Y^p = permanent income,

C^p = permanent consumption,

K= MPC

As per this theory,short run MPC is lower than long run MPC because a person is not sure that increase in income will retain in the long run or not.